The moon revolves round the earth 13 times in one year. if the ratio of the sun-earth distance to the earth-moon distance is 392, then the atio of the masses of sn and earthwill be............

Grade:12

2 Answers

Arun
6 years ago
Dear student
 
T²es= 4 (pi)² R³es/G Ms
T²me = 4 (pi)² R³me/G Me
[Tme/Tes]² =  R³me * Ms/ R³es *Me
(1/13)² = (1/392)³ Ms/Me
Ms/Me = 392³/13²
Regards
Arun (askIITians forum expert)
nirmal patel
6 years ago
the moon revolves round the earth 13 gtimes in one year if the ration of sun earth distance to earth moon distnece is 392 then the ratio of masses of sun and eaqrth will be